OpenStudy (anonymous):

Which of the following equations would graph a line parallel to 3y=2x-x+5 A) 3y=x+1 B) y=x+1 C) y=2/3x+1 D) 3/2y=x+1

OpenStudy (anonymous):

@RoseDryer

OpenStudy (luigi0210):

Combine like terms, ten put it in slope-intercept form

OpenStudy (anonymous):

to check if one line is parallel to the other they need to have the same slope

OpenStudy (anonymous):

I'm still confused..

OpenStudy (anonymous):

3y=2x-x+5 3y=x+5 Divide by 3 on both side y=1/3x+5/3 Find one that is parallel by finding one with the same slope. So get all into y=mx+b

OpenStudy (anonymous):

A) 3y=x+1 divide by 3 y=1/3x+1/3 B) y=x+1 C) y=2/3x+1 D) 3/2y=x+1 I would multiply by 2 on both sides 3y=2x+2 divide by 3 y=2/3x+2/3

OpenStudy (anonymous):

example y = 2x + 3 is parallel with y = 2x + 5 because the slope is the same , slope can be found by : m = y2 - y1 / x2 - x1 or by differenciation dy/dx or by just comparison with the general form y = mx + c where m is the slope hope that clears the confusion.

OpenStudy (anonymous):

okay so it would be A?

OpenStudy (anonymous):

 Yepp!!!!